he tens digit of a two-digit number is three greater than the ones digit. The difference between ... than the ones digit. The difference between twice the original number and half the number obtained after reversing the digit is 108.
Answer by lwsshak3(11628)

he tens digit of a two-digit number is three greater than the ones digit. The difference between ... than the ones digit. The difference between twice the original number and half the number obtained after reversing the digit is 108.
***
let u=units digit
let t=tens digit
t=u+3
2(10t+u)-(10u+t)/2=108
lcd:2
40t+4u-10u-t=216
40(u+3)+4u-10u-(u+3=216
40u+120+4u-10u-u-3=216
33u=99
u=3
t=u+3=6
original number: 63
check:
2*63-36/2=126-18=108